OpenStudy (anonymous):

19-13p=-17-5 Really unsure how to do these types of problems, any explanations would be greatly appreciated

OpenStudy (anonymous):

First, you have to isolate your variable.

OpenStudy (anonymous):

You do that by subtracting the 19 on both sides. Remember that what you do to one side you must always do to the other.

OpenStudy (anonymous):

so, now you have: -13p = -17-5-19

OpenStudy (anonymous):

So, when adding similar terms, you can combine, -17, -5, and -19 to get -41

OpenStudy (anonymous):

Now, you have -13p = -41p

OpenStudy (anonymous):

scratch that. I meant -13p = -41

OpenStudy (anonymous):

So, you divide by -13 to get the p alone.

OpenStudy (anonymous):

Your final answer: p = 41/13 = 3.15

OpenStudy (anonymous):

Okay?

OpenStudy (anonymous):

I think I understand it much better now. Thank you

OpenStudy (anonymous):

No problem! :)